OpenStudy (anonymous):

Write the following polynomial as products of linear factors. 9z^2-6z+5

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

it is not factorable :(

OpenStudy (anonymous):

The answer is (3z-1-2i)(3z-1+2i)... but how to get there?

OpenStudy (anonymous):

Try using the QUADRATIC FORMULA. :))

OpenStudy (anonymous):

Can u see if you get the answer?

OpenStudy (mani_jha):

Set the function equal to zero. Find the roots(using that discriminant formula). If the roots are a,b then the function can be rewritten as: \[(z-a)(z-b)\] I get\[(a,b)=(6\pm \sqrt{36-180})/18\]

OpenStudy (anonymous):

Which comes out as\[ 1\pm2i \over 3\]

OpenStudy (anonymous):

How does 3 become 3z and 1 become -1?

OpenStudy (mani_jha):

\[a=(1+2i)/3; b=(1-2i)/3\] \[(z-(1+2i)/3)(z-(1-2i)/3)=0\] Multiply this by 3

OpenStudy (anonymous):

Thanks :)